OpenStudy (anonymous):

Use an angle sum identity to verify the identity cos2x=2cos^2x-1

jimthompson5910 (jim_thompson5910):

cos(2x) = 2*cos^2(x) - 1 cos(x+x) = 2*cos^2(x) - 1 cos(x)cos(x) - sin(x)sin(x) = 2*cos^2(x) - 1 cos^2(x) - sin^2(x) = 2*cos^2(x) - 1 cos^2(x) - (1 - cos^2(x)) = 2*cos^2(x) - 1 cos^2(x) - 1 + cos^2(x) = 2*cos^2(x) - 1 2cos^2(x) - 1 = 2*cos^2(x) - 1

jimthompson5910 (jim_thompson5910):

The idea is to transform one side into the other and you can only work on ONE side (leaving the other unchanged)

OpenStudy (anonymous):

ok thank you I was always trying to work out both sides.

jimthompson5910 (jim_thompson5910):

yw